# Zeolite Adsorption Property Prediction using Deep Learning

### v1

@@ Abstract
The ability to efficiently predict adsorption properties of zeolites can be of large benefit in accelerating the design process of novel materials. The existing configuration space for these materials is wide, while existing molecular simulation methods are computationally expensive. In this work, we propose a model which is 4 to 5 orders of magnitude faster at adsorption properties compared to molecular simulations. To validate the model, we generated datasets containing various aluminium configurations for the MOR, MFI, RHO and ITW zeolites along with their heat of adsorptions and Henry coefficients for CO, obtained from Monte Carlo simulations. The predictions obtained from the Machine Learning model are in agreement with the values obtained from the Monte Carlo simulations, confirming that the model can be used for property prediction. Furthermore, we show that the model can be used for identifying adsorption sites. Finally, we evaluate the capability of our model for generating novel zeolite configurations by using it in combination with a genetic algorithm.
@@ 1 Introduction
Over the past years, the amount of CO in the atmosphere has been increasing, and greenhouse effects have become increasingly evident. A potential way to reduce the carbon levels in the atmosphere is carbon capture 1. Nanoporous materials, such as zeolites, are good candidates 2, due to their high adsorption capacity of gasses 3, 4, 5. In addition, zeolites have a high thermal stability 6, and their synthesis can have a low cost 7 compared to other adsorbents. Furthermore, there exists a large number of synthesizable zeolite topologies 8, with different pore sizes and properties. For each zeolite topology, there exist multiple possible configurations depending on the amount of aluminium and silicon atoms. Here, the Si/Al ratio can influence the properties of the zeolite 9, 10, 11, such as the adsorption of CO. While the overall trend shows that decreasing the Si/Al ratio leads to a higher heat of adsorption, there can still be a lot of variance in the heat of adsorption for a single framework with a set Si/Al ratio 12.
@@
In this work, we extend the EPCN model 35 in such that it allows us to identify the adsorption sites, without a significant change in performance. This makes the model interpretable, as we can now identify the pores responsible for the adsorption. Furthermore, we extend the dataset provided with ECPN (CO heat of adsorption for MOR and MFI) with CO heat of adsorption values for different configurations of the RHO and ITW zeolites, as well as with Henry coefficients for all four zeolites. In addition, we extend the model to predict the Henry coefficient of each zeolite configuration in the dataset. Using this dataset, we show that the model performs excellently predicting the CO heat of adsorption and Henry coefficient simultaneously, while being significantly faster than Monte Carlo simulations. Finally, we demonstrate that this model can be used for inverse design of zeolites based on their heat of adsorption, by using the model for the fitness function of a genetic algorithm.
@@ 2.2 Computational Details
In this study, we investigated the heat of adsorption () and Henry coefficient () 41, 42, which reflect the interaction strength between the CO and the zeolite. To calculate these two properties, Monte Carlo (MC) simulations using the Widom particle insertion method in the canonical ensemble (NVT) were performed 43. To obtain the heat of adsorption value, Equation 1 was used, where represents the internal energy difference before and after the adsorption of the guest molecule. In this equation, we neglect the contributions to the internal energy of the zeolite and the CO molecules, since they are modeled as rigid bodies.
@@ 3.2 Interpretability
As a result of our proposed change in architecture, the model now predicts a scalar value per feature for each pore. In turn, the predictions can be interpreted as the contribution of each pore to the heat of adsorption/Henry coefficient. To qualitatively verify whether this is the case, we sampled 1500 CO positions from the MC simulation for 4 structures. In Figure 6, we visualised the CO distribution and the final pore values for these zeolites. It can be seen that the final values of the pore embeddings are proportional to the amount of times there was a CO molecule present in the pore. For example, in the top left and bottom right cases, the CO molecule is mainly present in the outer 12 ring. This is reflected in the final pore values, which in both cases is higher for the outer 12 ring. As such, the output of the model can be used to estimate which pores are responsible for the CO adsorption.
@@ 4 Conclusions
Overall, we have shown that ML can be a useful tool for modelling zeolites, with our proposed method achieving excellent performance on the heat of adsorption and Henry coefficients predictions, while also showing that this class of models could be interpretable. Our model can provide useful insights into the underlying mechanism that triggers CO adsorption, such as the the contribution of each framework region into the heat of adsorption value. This strategy opens the possibility of identifying adsorbent features to maximize the performance of a target property. In addition, we have demonstrated that these models are a step towards inverse design of new materials, as it is possible to optimize structures for a target heat of adsorption using our model in combination with a genetic algorithm.

### v2

# Graph Neural Networks for Carbon Dioxide Adsorption Prediction in Aluminium-Exchanged Zeolites

@@ Abstract
The ability to efficiently predict adsorption properties of zeolites can be of large benefit in accelerating the design process of novel materials. The existing configuration space for these materials is wide, while existing molecular simulation methods are computationally expensive. In this work, we propose a model which is 4 to 5 orders of magnitude faster at adsorption properties compared to molecular simulations. To validate the model, we generated datasets containing various aluminium configurations for the MOR, MFI, RHO and ITW zeolites along with their heat of adsorptions and Henry coefficients for CO2, obtained from Monte Carlo simulations. The predictions obtained from the Machine Learning model are in agreement with the values obtained from the Monte Carlo simulations, confirming that the model can be used for property prediction. Furthermore, we show that the model can be used for identifying adsorption sites. Finally, we evaluate the capability of our model for generating novel zeolite configurations by using it in combination with a genetic algorithm.
@@ 1 Introduction
Over the past years, the amount of CO2 in the atmosphere has been increasing, and greenhouse effects have become increasingly evident. A potential way to reduce the carbon levels in the atmosphere is carbon capture 1. Nanoporous materials, such as zeolites, are good candidates 2, due to their high adsorption capacity of gasses 3, 4, 5. In addition, zeolites have a high thermal stability 6, and their synthesis can have a low cost 7 compared to other adsorbents. Furthermore, there exists a large number of synthesizable zeolite topologies 8, with different pore sizes and properties. For each zeolite topology, there exist multiple possible configurations depending on the amount of aluminium and silicon atoms. Here, the Si/Al ratio can influence the properties of the zeolite 9, 10, 11, such as the adsorption of CO2. While the overall trend shows that decreasing the Si/Al ratio leads to a higher heat of adsorption, there can still be a lot of variance in the heat of adsorption for a single framework with a set Si/Al ratio 12.
@@
In this work, we extend the EPCN model 35 in such that it allows us to identify the adsorption sites, without a significant change in performance. This makes the model interpretable, as we can now identify the pores responsible for the adsorption. Furthermore, we extend the dataset provided with ECPN (CO2 heat of adsorption for MOR and MFI) with CO2 heat of adsorption values for different configurations of the RHO and ITW zeolites, as well as with Henry coefficients for all four zeolites. In addition, we extend the model to predict the Henry coefficient of each zeolite configuration in the dataset. Using this dataset, we show that the model performs excellently predicting the CO2 heat of adsorption and Henry coefficient simultaneously, while being significantly faster than Monte Carlo simulations. Finally, we demonstrate that this model can be used for inverse design of zeolites based on their heat of adsorption, by using the model for the fitness function of a genetic algorithm.
@@ 2.2 Computational Details
In this study, we investigated the heat of adsorption () and Henry coefficient () 41, 42, which reflect the interaction strength between the CO2 and the zeolite. To calculate these two properties, Monte Carlo (MC) simulations using the Widom particle insertion method in the canonical ensemble (NVT) were performed 43. To obtain the heat of adsorption value, Equation 1 was used, where represents the internal energy difference before and after the adsorption of the guest molecule. In this equation, we neglect the contributions to the internal energy of the zeolite and the CO2 molecules, since they are modeled as rigid bodies.
@@ 3.2 Interpretability
As a result of our proposed change in architecture, the model now predicts a scalar value per feature for each pore. In turn, the predictions can be interpreted as the contribution of each pore to the heat of adsorption/Henry coefficient. To qualitatively verify whether this is the case, we sampled 1500 CO2 positions from the MC simulation for 4 structures. In Figure 6, we visualised the CO2 distribution and the final pore values for these zeolites. It can be seen that the final values of the pore embeddings are proportional to the amount of times there was a CO2 molecule present in the pore. For example, in the top left and bottom right cases, the CO2 molecule is mainly present in the outer 12 ring. This is reflected in the final pore values, which in both cases is higher for the outer 12 ring. As such, the output of the model can be used to estimate which pores are responsible for the CO2 adsorption.
@@ 4 Conclusions
Overall, we have shown that ML can be a useful tool for modelling zeolites, with our proposed method achieving excellent performance on the heat of adsorption and Henry coefficients predictions, while also showing that this class of models could be interpretable. Our model can provide useful insights into the underlying mechanism that triggers CO2 adsorption, such as the the contribution of each framework region into the heat of adsorption value. This strategy opens the possibility of identifying adsorbent features to maximize the performance of a target property. In addition, we have demonstrated that these models are a step towards inverse design of new materials, as it is possible to optimize structures for a target heat of adsorption using our model in combination with a genetic algorithm.